OpenStudy (amina_soneviseth):

8 ( y - 5 ) + 2y = (-10)

OpenStudy (anonymous):

1.25

OpenStudy (amina_soneviseth):

@CJDOG how do i solve it?

OpenStudy (anonymous):

8 ( y - 5 ) + 2y = (-10) ÷8 ÷8 y-5 + 2y = -1.25 +5 +5 y+2y = 3.75 3y = 3.75 ÷3 ÷3 y=1.25

OpenStudy (anonymous):

hows that?

OpenStudy (anonymous):

@amina_soneviseth

OpenStudy (amina_soneviseth):

@CJDOG thank you, i only asked because the rule is NO direct answers